Grandma Betty is between fifty and seventy years old. Each of her sons has as many sons as brothers. The total number of Betty’s sons and grandsons is equal to the number of years that are in her age. How many years old is Betty?
Answer by ankor@dixie-net.com(22740)

Grandma Betty is between fifty and seventy years old.
Each of her sons has as many sons as brothers.
The total number of Betty’s sons and grandsons is equal to the number of years that are in her age.
How many years old is Betty?
:
let s = no. of son B has
then
(s-1) = no. of sons her sons have
:
s(s-1) = 50 - 70
we know that the 2 factors of her age differ by 1, that would be 8*7 = 56
s(s-1) = 56
s^2 - s - 56 = 0
Factors to
(s-8)(s+7) = 0
positive solution
s = 8 sons
Therefore she has 8 sons, with 7 sons of their own
8*7 = 56 is B's age